Interpret the initial value problem
0
0
0
2
2
2
0, (0) , w
q
b q q q
q
= 

+ = =
t= dt
d
dt
d
for any physical situation and hence solve the problem
1
Expert's answer
2020-03-23T12:40:54-0400

Interpret the initial value problem


"{d^2 \theta \over dt^2}+\beta^2\theta=0, \ \theta(0)=\theta_0,\ {d\theta \over dt}(0)=\omega_0"

Use the notation


"{d^2 \theta \over dt^2}=\theta'',\ {d \theta \over dt}=\theta'"

We get


"\theta''+\beta^2\theta=0"

The auxiliary equation is "r^2+\beta^2=0" or "r^2=-\beta^2," whose roots are "\pm i\beta."

The general solution is


"\theta(t)=c_1 \cos(\beta t)+c_2\sin(\beta t)"

Then


"\theta'=-\beta c_1\sin(\beta t)+\beta c_2\cos(\beta t)"

The initial conditions become


"\theta(0)=c_1=\theta_0"

"\theta'(0)=\beta c_2=\omega_0=>c_2={\omega_0 \over \beta}"

Hence


"\theta(t)=\theta_0 \cos(\beta t)+{\omega_0 \over \beta}\sin(\beta t)"


Spring system


"\begin{matrix}
 x & displacement \\
 dx/dt & velocity \\
 m & mass \\
 c & damping\ constant \\
 k & spring\ constant \\
 F(t) & external\ force
\end{matrix}"

If we ignore any external resisting forces (due to air resistance or friction) then, by Newton’s Second Law (force equals mass times acceleration), we have


"m{d^2 x \over dt^2}=-kx"

Or


"{d^2 x \over dt^2}+{k\over m}x=0"

Let "\beta^2=\dfrac{k}{m}." Then the solution of the initial value problem


"{d^2 x \over dt^2}+\beta^2x=0, \ x(0)=x_0,\ {dx\over dt}(0)=v_0"

is


"x(t)=x_0 \cos(\beta t)+{v_0 \over \beta}\sin(\beta t), \beta=\sqrt{{k \over m}}"